Ratio of maximum to minimum value of the expression 1 - 4sintheta + 3cos^2theta isa)-16/3b)-16/7c)-16/9d)None of thesePls. give detailed answer

Grade:10

1 Answers

Harsha
6 years ago
Denote cosx as c,sin x as s.......Y = 1-4s+3c^2=1-4s+3(1-s^2)=4-4s-3s^2......differentiate this equation.....dy/dthetha = -4c-6cs=0......c=0,s=-2/3......by substituting this in y we get maximum and minimum as 14/3,-3....the ratio of maximum to minimum is -14/9....therefore d)none of these is must be the answer